The Magellanic Clouds Newsletter

An electronic exchange on Magellanic Clouds research

Edited by Eva K. Grebel and You-Hua Chu

         mcnews@astro.uiuc.edu
http://www.astro.uiuc.edu/mcnews/MCNews.html
 http://www.astro.uni-bonn.de/~mcnews/

No. 28 January 10, 1999

Contents

News  1
Abstracts of 11 refereed papers 2
Job opportunity  9

News

Catalog of Eclipsing Binary Stars in the SMC available over the Internet

The OGLE-II microlensing search team announces release of the catalog of eclipsing binary stars from central part of the Small Magellanic Cloud. The catalog contains 1459 eclipsing systems.

The catalog, finding charts and BVI photometry data for all objects are available from the OGLE Internet archive:

http://www.astrouw.edu.pl/~ftp/ogle or its mirror
http://www.astro.princeton.edu/~ogle

Andrzej Udalski

Best wishes for a happy and productive year 1999!

Abstracts of Refereed Papers

The Supergiant Shell LMC 2: I. The Kinematics and Physical Structure

S. D. Points^{1,6}, Y.-H. Chu^{1,6}, S. Kim^{1,2}, R. C. Smith^{3,6}, S. L. Snowden^{4,7}, W. Brandner^{1,8}, R. A. Gruendl^{5}

1 Astronomy Dept., University of Illinois, 1002 W. Green St. Urbana, IL 61801, USA 2 Mount Stromlo and Siding Spring Observatories, Weston Creek Post Office, ACT 2611, Australia 3 Dept. of Astronomy, University of Michigan, 834 Dennison, Ann Arbor, MI 48109, USA 4 NASA/Goddard Space Flight Center, Code 662, Greenbelt, MD 20771, USA 5 Laboratory for Astron. Imaging, Astronomy Dept., University of Illinois, 1002 W. Green St., Urbana, IL 61801, USA 6 Visiting Astronomer, Cerro Tololo Inter-American Observatory (CTIO), National Optical Astronomy Obs. (NOAO) 7 Universities Space Research Association 8 Caltech - JPL/IPAC, Mail Code 100-22, Pasadena, CA, 91125, USA

LMC 2 has the brightest most coherent filamentary structure of all known supergiant shells in the Large Magellanic Cloud. The optical emission line images show active star formation regions along the western edge and long filaments to the east. ROSAT PSPC and HRI images show bright Xray emission from within the shell boundary, indicating the presence of hot gas. Counter-intuitively, neither high-resolution echelle spectra in the Hα line nor aperture synthesis H i 21-cm emission line observations show LMC 2 to have the kinematics expected of an expanding shell. Rather, LMC 2 appears to consist of hot gas confined between H i sheets. The interior surfaces of these sheets are ionized by the UV flux of massive stars in the star formation regions along the periphery of LMC 2, while the heating is provided by outflows of hot gas from the star formation regions and by SNRs interior to LMC 2.

We have compared LMC 2 to other supergiant shells in the LMC and in more distant galaxies. When the spatial resolution of our data are degraded, we find that LMC 2 resembles supergiant shells observed at a distance of 4 Mpc that have previously been interpreted as expanding shells. Therefore, great caution should be exercised in the analysis and interpretation of the kinematics of distant supergiant shells to prevent overestimates of their velocities and total kinetic energies.

Accepted by: The Astrophysical Journal

For preprints, contact points@astro.uiuc.edu
Also available from the URL http://www.astro.uiuc.edu/~points

On the model of dust in the Small Magellanic Cloud

Victor G. Zubko^{1,2}

1 Department of Physics, Technion – Israel Institute of Technology, Haifa 32000, Israel 2 On leave from the Main Astronomical Observatory, National Academy of Sciences of the Ukraine, Kiev

We present here dust models for the Small Magellanic Cloud bar calculated for the first time with the regularization approach. A simple mixture of the core–mantle and/or composite grains (mostly made from silicates and organic refractory) together with silicon nanoparticles is consistent with the following: 1) the observed extinction toward AzV 398 (a typical SMC bar sightline), 2) the elemental abundances, and 3) the strength of the interstellar radiation field. We predict the expected albedo and asymmetry parameter of the models, which can be tested in future observations. The proposed dust models can also be tested by looking for the expected extended red emission.

Accepted by: Astrophysical Journal Letters

For preprints, contact zubko@phquasar.technion.ac.il
Also available from the URL http://physics.technion.ac.il/~zubko/eb.html

The detection of extragalactic ^{15}N:

Consequences for nitrogen nucleosynthesis and chemical evolution

Y.-N. Chin^{1}, C. Henkel^{2,3}, N. Langer^{4}, R. Mauersberger^{5}

1 Institute of Astronomy and Astrophysics, Academica Sinica, P.O.Box 1-87 Nankang, 11529 Taipei, Taiwan 2 Max-Planck-Institut für Radioastronomie, Auf dem Hügel 69, D-53121 Bonn, Germany 3 European Southern Observatory, Casilla 19001, Santiago 19, Chile 4 Institut für Physik, Universität Potsdam, Am Neuen Palais 10, D-14469 Potsdam, Germany 5 Steward Observatory, The University of Arizona, Tucson, AZ 85721, USA

Detections of extragalactic ^{15}N are reported from observations of the rare hydrogen cyanide isotope HC^{15}N toward the Large Magellanic Cloud (LMC) and the core of the (post-) starburst galaxy NGC 4945. Accounting for optical depth effects, the LMC data from the massive star-forming region N113 infer a ^{14}N/^{15}N ratio of 111 ± 17, about twice the ^{12}C/^{13}C value. For the LMC star-forming region N159HW and for the central region of NGC 4945, ^{14}N/^{15}N ratios are also ≈ 100. The ^{14}N/^{15}N ratios are smaller than all interstellar nitrogen isotope ratios measured in the disk and center of the Milky Way, strongly supporting the idea that ^{15}N is predominantly of ‘primary’ nature, with massive stars being its dominant source. Although this appears to be in contradiction with standard stellar evolution and nucleosynthesis calculations, it supports recent findings of abundant ^{15}N production due to rotationally induced mixing of protons into the helium-burning shells of massive stars.

Accepted by: Astrophysical Journal Letters

For preprints, contact einmann@asiaa.sinica.edu.tw
Also available from the URL http://xxx.lanl.gov/abs/astro-ph/9812375

Dust in R 71: First detection of crystalline silicates in the LMC

R.H.M. Voors^{1,2}, L.B.F.M. Waters^{3,4}, P.W. Morris^{1,3}, N.R. Trams^{5},

A. de Koter^{3} and J. Bouwman^{3}

1 SRON Laboratory for Space Research, Sorbonnelaan 2, 3584 CA Utrecht, The Netherlands 2 Astronomical Institute, University of Utrecht, Princetonplein 5, 3508 TA Utrecht, The Netherlands 3 Astronomical Institute Anton Pannekoek, University of Amsterdam, Kruislaan 403, 1098 SJ Amsterdam, The Netherlands 4 SRON Laboratory for Space Research, PO Box 800, 9700 AV Groningen, The Netherlands 5 Integral Science Operations, Astrophysics Division of ESA, ESTEC SCI-SAG, PO Box 299, 2200 AG Noordwijk, The Netherlands

We present infrared spectroscopy taken with the Infrared Space Observatory (ISO) of the Luminous Blue Variable (LBV) R71 in the Large Magellanic Cloud (LMC). The spectrum shows clear evidence for the presence of crystalline olivine at 23.5 μm. This is the first detection of circumstellar crystalline silicates outside our galaxy. In addition, we identify emission at 6.2, 7.7 and possibly 8.6 μm from C-rich small grains (PAHs). The presence of C-rich grains is not expected in an environment where C/O is less than 1. We fit the dust spectrum using a radiative transfer model and find a dust mass of 0.02 M_{⊙}. R71 was probably a Red Supergiant when it produced the dust shell and had a time-averaged mass loss rate of the order of 7 × 10^{−4} M_{⊙} yr^{−1} for a gas/dust ratio of 100.

To appear in: Astronomy and Astrophysics, 341, L67 For preprints, contact voors@phys.uu.nl

Also available from the URL http://www.phys.uu.nl/ voors/papers.html
or by anonymous ftp at ruunat.phys.uu.nl/pub/astronomy/voors/r71/r71.ps

The Circumstellar Extinction of Planetary Nebulae

R. Ciardullo^{1} and G.H. Jacoby^{2}

1Department of Astronomy and Astrophysics, Penn State University, 525 Davey Lab, University Park, PA 16802 2Kitt Peak National Observatory, National Optical Astronomy Observatories, P.O. Box 26732, Tucson, AZ 85726

We analyze the dependence of circumstellar extinction on core mass for the brightest planetary nebulae (PNe) in the Magellanic Clouds and M31. We show that in all three galaxies, a statistically significant correlation exists between the two quantities, such that high core mass objects have greater extinction. We model this behavior, and show that the relation is a simple consequence of the greater mass loss and faster evolution times of high mass stars. The relation is important because it provides a natural explanation for the invariance of the [O III] λ5007 planetary nebula luminosity function (PNLF) with population age: bright Population I PNe are extinguished below the cutoff of the PNLF. It also explains the counter-intuitive observation that intrinsically luminous Population I PNe often appear fainter than PNe from older, low-mass progenitors.

Accepted by: The Astrophysical Journal For preprints, contact rbc@astro.psu.edu

Also available from the URL http://www.astro.psu.edu/users/rbc/chem pubs.html

Spatially Resolved Nebulae around the Ofpe/WN9 stars S61 and BE381

A. Pasquali^{1}, A. Nota^{2} and M. Clampin^{2}

1 ST-ECF/ESO, Karl-Schwarzschild-Strasse 2, D-85748, Garching bei München, Germany 2 Space Telescope Science Institute, 3700 San Martin Drive, Baltimore, MD 21218, USA

We present new high-resolution coronographic imaging and medium-resolution spectroscopy of the circumstellar region around S61 and BE381, two Ofpe/WN9 stars in the Large Magellanic Cloud. The observations were carried out at the ESO/NTT (La Silla) in January 1996 and July 1998. The excellent seeing conditions allowed the circumstellar nebulae associated with both S61 and BE381 to be spatially resolved for the first time. The nebula surrounding S61 has the appearance of a shell with a mild central axisymmetry. The surface brightness is not uniform, and the northern region of the nebula is the brightest. The nebula exhibits a bipolar structure with an overall morphology very similar to nebulae around other LBVs or Opfe/WN9 stars, especially S119. The diameter of the shell is 7.3′′, corresponding to a linear size of 1.8 pc. From the profile of nebular emission lines we clearly detect an expansion motion with a velocity of 28 km s^{−1}, which indicates a dynamical age of ∼ 30000 yrs. We find an electron density of 400 cm^{−3} and an electron temperature of 6120 K. The nebula is simular to other LBV nebulae in that it is nitrogen enriched. The observed chemical and dynamical properties confirm that the nebula is associated with the central star and is of stellar origin. This result implies that S61 is likely to have undergone a LBV-type outburst and, therefore, strengthens the suggestion that Ofpe/WN9 stars are quiescent LBVs. The situation is different for BE381. The Hα images of BE381 also reveal the presence of a faint nebulosity around the star; most of the nebular flux appears to be emitted by an arc of gas located to the east of BE381, while a much dimmer arc is detectable on the western side. The arcs delineate a shell of 13′′ in diameter, corresponding to a linear size of 3.2 pc, which appears to be expanding with a velocity of 14 km s^{−1}. From the nebular emission lines we derive an electron density ranging between 30 cm^{−3} and 120 cm^{−3} (assuming T_{e} = 10000 K), and a N^{+}/S^{+} ratio between 1.5 and 2.3, which are typical of HII regions. We therefore conclude that the shell detected around BE381 is not of stellar origin and probably represents the relic of the interstellar bubble blown by BE381 during its O main-sequence phase.

Accepted by: Astronomy & Astrophysics For preprints, contact apasqual@eso.org

A ROSAT PSPC X-Ray Survey of the Small Magellanic Cloud

P. Kahabka ^{1} W. Pietsch ^{2} M.D. Filipović ^{2,3,4} and F. Haberl^{2}

1 Astronomical Institute and Center for High Energy Astrophysics, University of Amsterdam, Kruislaan 403, NL–1098 SJ Amsterdam, The Netherlands 2 Max-Planck-Institut für extraterrestrische Physik, D–85740 Garching bei München, Germany 3 University of Western Sydney, Nepean, P.O. Box 10, Kingswood, NSW 2747, Australia 4 Australia Telescope National Facility, CSIRO, P.O. Box 76, Epping, NSW 2121, Australia

We present the results of a systematic search for point-like and moderately extended soft (0.1-2.4 keV) X-ray sources in a raster of nine pointings covering a field of 8.95 deg^{2} and performed with the ROSAT PSPC between October 1991 and October 1993 in the direction of the Small Magellanic Cloud (SMC). We detect 248 objects which we include in the first version of our SMC catalogue of soft X-ray sources. We set up seven source classes defined by selections in the count rate, hardness ratio and source extent. We find five high luminosity super-soft sources (1E 0035.4-7230, 1E 0056.8-7146, RX J0048.4-7332, RX J0058.6-7146 and RX J0103-7254), one low-luminosity super-soft source RX J0059.6-7138 correlating with the planetary nebula L357, 51 candidate hard X-ray binaries including eight bright hard X-ray binary candidates, 19 supernova remnants (SNRs), 19 candidate foreground stars and 53 candidate background active galactic nuclei (and quasars). We give a likely classification for ∼60% of the catalogued sources. The total count rate of the detected point-like and moderately extended sources in our catalogue is 6.9±0.3 s^{−1}, comparable to the background subtracted total rate from the integrated field of ∼ 6.1 ± 0.1 s^{−1}.

Accepted by: Astronomy & Astrophysics Supplements For preprints, contact ptk@astro.uva.nl

Also available from the URL xxx.lanl.gov/abs/astro-ph/9812143

The Optical Gravitational Lensing Experiment. Eclipsing Binary Stars in the Small Magellanic Cloud

A. Udalski^{1}, I. Soszyński^{1}, M. Szymański^{1} M. Kubiak^{1}, G. Pietrzyński^{1}, P. Woźniak^{2}, and K. Żebruń^{1}

1 Warsaw University Observatory, Al. Ujazdowskie 4, 00-478 Warszawa, Poland 2 Princeton University Observatory, Princeton, NJ 08544-1001, USA

We present the catalog of 1459 eclipsing binary stars detected in the central 2.4 square degree area of the Small Magellanic Cloud during the OGLE-II microlensing search. The sample includes objects brighter than I ≈ 20 mag with periods ranging from about 0.3 to 250 days. The average completeness of the catalog is about 80%. Statistics of the entire sample and well detached systems, suitable for distance determination, are also presented.

The catalog, finding charts and BVI photometry data for all objects are available from the OGLE Internet archive.

Accepted by: Acta Astronomica 48, 563.

For preprints, contact udalski@sirius.astrouw.edu.pl
Also available from the URL http://xxx.lanl.gov/abs/astro-ph/9812348

Interacting star clusters in the LMC Overmerging problem solved by cluster group formation

Stéphane Leon^{1,2}, Gilles Bergond^{2,3}, and Antonella Vallenari^{4}

1 DEMIRM, Observatoire de Paris, 61, Av. de l’Observatoire, F-75014 Paris, France 2 CAI-MAMA, Observatoire de Paris, 61, Av. de l’Observatoire, F-75014 Paris, France 3 DASGAL, Observatoire de Paris-Meudon, 5, Place J. Janssen F-92195 Meudon, France 4 Astronomical Observatory of Padova, Vicolo dell’Osservatorio 5, I-35122 Padova, Italy

We present the tidal tail distributions of a sample of candidate binary clusters located in the bar of the Large Magellanic Cloud (LMC). One isolated cluster, SL 268, is presented in order to study the effect of the LMC tidal field. All the candidate binary clusters show tidal tails, confirming that the pairs are formed by physically linked objects. The stellar mass in the tails covers a large range, from 1.8 × 10^{3} to 3 × 10^{4} M_{⊙}. We derive a total mass estimate for SL 268 and SL 356. At large radii, the projected density profiles of SL 268 and SL 356 fall off as r^{−γ}, with γ = 2.27 and γ = 3.44, respectively. Out of 4 pairs or multiple systems, 2 are older than the theoretical survival time of binary clusters (going from a few 10^{6} years to 10^{8} years). A pair shows too large age difference between the components to be consistent with classical theoretical models of binary cluster formation (Fujimoto & Kumai 1997).

We refer to this as the “overmerging” problem. A different scenario is proposed: the formation proceeds in large molecular complexes giving birth to groups of clusters over a few 10^{7} years. In these groups the expected cluster encounter rate is larger, and tidal capture has higher probability. Cluster pairs are not born together through the splitting of the parent cloud, but formed later by tidal capture. For 3 pairs, we tentatively identify the star cluster group (SCG) memberships. The SCG formation, through the recent cluster starburst triggered by the LMC-SMC encounter, in contrast with the quiescent open cluster formation in the Milky Way can be an explanation to the paucity of binary clusters observed in our Galaxy.

Accepted by: Astronomy & Astrophysics Main Journal

For preprints, contact Gilles.Bergond@obspm.fr
Also available from the URL http://xxx.lanl.gov/astro-ph/9812112

Microlensing towards the Small Magellanic Cloud — EROS 2 two-year analysis

EROS Collaboration

We present here the analysis of the first two years of data towards the Small Magellanic Cloud with the EROS (Expérience de Recherche d’Objets Sombres) project. A single event is detected, already present in the first year analysis. This low event rate allows us to constrain the halo fraction in the mass range [10^{−2}, 1] M_{⊙}.

Submitted to: Astrophysical Journal Letters

For preprints, contact aubourg@hep.saclay.cea.fr
Also available from the URL http://xxx.lanl.gov/abs/astro-ph/9812173

Observations of the Binary Microlens Event MACHO-98-SMC-1 by the Microlensing Planet Search Collaboration

S.H. Rhie^{1}, A.C. Becker^{1,2}, D.P. Bennett^{1,3}, P.C. Fragile^{1}, B.R. Johnson^{5}, L.J. King^{1,3}, B.A. Peterson^{4}, and J. Quinn^{1}

(The Microlensing Planet Search Collaboration)

1 Department of Physics, University of Notre Dame, Notre Dame, IN 46556, USA 2 Departments of Astronomy and Physics, University of Washington, Seattle, WA 98195, USA 3 Center for Particle Astrophysics, University of California, Berkeley, CA 94720, USA 4 Mt. Stromlo and Siding Spring Observatories, Australian National University, Weston, ACT 2611, Australia 5 Tate Laboratory of Physics, University of Minnesota, Minneapolis, MN 55455, USA

We present the observations of the binary lensing event MACHO-98-SMC-1 conducted at the Mt. Stromlo 74” telescope by the Microlensing Planet Search (MPS) collaboration. The MPS data constrain the first caustic crossing to have occurred after 1998 June 5.55 UT and thus directly rule out one of the two fits presented by the PLANET collaboration (model II). This substantially reduces the uncertainty in the the relative proper motion estimations of the lens object.

We perform joint binary microlensing fits of the MPS data together with the publicly available data from the EROS, MACHO/GMAN and OGLE collaborations. We also study the binary lens fit parameters previously published by the PLANET and MACHO/GMAN collaborations by using them as initial values for χ^{2} minimization. Fits based on the PLANET model I appear to be in conflict with the GMAN-CTIO data. From our best fit, we find that the lens system has a proper motion of μ = 1.3 ± 0.2 km s^{−1} kpc^{−1} with respect to the source, which implies that the lens system is most likely to be located in the Small Magellanic Cloud in agreement with previous reports.

Submitted to: The Astrophysical Journal

Also available from the URL http://xxx.lanl.gov/abs/astro-ph/9812252

Job Opportunity

Postdoctoral Position at University of Arizona, Steward Observatory

The University of Arizona Department of Astronomy is soliciting applications for postdoctoral positions working with Dr. Dennis Zaritsky beginning fall 1999 at Steward Observatory. One, and possibly two, positions will be filled. The target research areas include the properties of dust in galaxies, the stellar populations of the Magellanic Clouds, and the properties of high redshift galaxy clusters. Within this range of topics, the successful applicant(s) will define their research to best match their interests. The successful applicant will have access to Dr. Zaritsky’s ongoing Magellanic Clouds Photometric Survey and galaxy cluster survey, and to the full range of facilities at Steward Observatory (which include the 6.5m MMT, the 2.3 Bok Telescope, the SMT, and several smaller telescopes: see http://www.as.arizona.edu).

Applicants must be able to assume the position in the fall of 1999 and must complete all requirements for a Ph.D. in astronomy or related field before assuming the position. The position is for two years, with the possibility of a merit-based extension for a third year. Applicants should send a curriculum vitae, bibliography, a statement of research interests, and arrange for three reference letters to be sent to

Prof. Dennis Zaritsky

Steward Observatory

University of Arizona

Tucson, Arizona 85721

USA

Applicant review will begin February 1, 1999, and continue until at least one position is filled. Please refer to job # 13012.

The University of Arizona is an EEO/AA employer M/W/D/V.